the perimeter of the rectangle is 50 write a linear equation in two variables, to find x when y is 20
Answer by racoon(5)

In a rectangle, there are two pairs of parallel lines.
If we labelled a length of one pair of lines as x and the other length pair of lines as y, the equation will look like this;
2x + 2y = 50
Subtitutes y = 20
2x + 2(20) = 50
2x + 40 = 50
2x = 50 - 40
2x = 10
x = 10/2
x = 5